# AI-Driven Robotics for Free-Space Optics

### v1

@@ X.3 Validation of optical setups
To ensure that optical setups generated by the LLM are accurate and adhere to physical and geometric constraints, we employ a validation procedure that extends the checks used during synthetic dataset generation. First, we apply a set of rule-based spatial constraints to the model-generated setup: all components must lie within the robot’s 61 cm radius reach, remain outside the 15 cm radius circular footprint of the robot at the origin, maintain non-overlapping placement based on their 7.5 cm × 6.4 cm physical dimensions, and internal beam paths must not intersect the robotic arm’s footprint.
@@
Passing these spatial checks does not guarantee that the laser beam will correctly propagate through the setup, as the component orientations or positions may still be physically inaccurate. To address this, we perform an additional validation step that verifies the structural correctness of the generated setup against a known reference. We do this by attempting to identify a similarity transformation that maps the LLM-generated setup onto one of the predefined reference setup of the same type. This is framed as a geometric alignment task, where the candidate transformation must preserve component identities and maintain approximate spatial relationships within tolerance bounds. If such a transformation exists, the generated setup is deemed valid due to its equivalence to the corresponding reference configuration. This step ensures that the model has not only obeyed spatial constraints but also produced a layout that structurally corresponds to an accurate optical arrangement of components. Furthermore, this allows us to benchmark and test the accuracy of LLM-generated setups with various prompting strategies (Figure 2b of the main text).
@@ X.4 Generation of robot motion code
[…]
X.5 Computer vision system
@@
The robotic system’s vision pipeline is built around a stereo vision setup composed of two 4K ELP cameras mounted 23 cm apart. Initial calibration is conducted using a standard checkerboard pattern, with images collected at varying positions on the optical workbench. Intrinsic and extrinsic parameters of the stereo pair are obtained using the OpenCV cv2.calibrateCamera and cv2.stereoCalibrate functions. The calibration outputs are validated by comparing the predicted baseline offset to the physical camera separation, confirming sub-millimeter discrepancies. These calibration parameters are stored in a serialized NumPy archive (stereo_calibration2.npz) for use in subsequent 3D position estimations during coarse alignment operations.
@@
To enable robust and accurate object detection, we employ the ArUco fiducial markers. Each optical component is tagged with a unique identifier drawn from a 6×6 binary dictionary comprising over 800 distinguishable codes. These tags are detected using the cv2.aruco module, which computes the component’s pose via the cv2.solvePnP algorithm using previously derived camera intrinsics. While the x and y positional estimates from ArUco detection are sufficiently accurate for coarse positioning, the z-coordinate is instead inferred from a depth measurement. For this, we deploy a robot-mounted Intel RealSense depth camera equipped with LiDAR. After the ArUco tag is centered in the camera field of view, the LiDAR sensor provides a high-accuracy depth estimate, enabling precise z-coordinate acquisition. The robot’s end-effector is then translated vertically from a known camera-to-gripper offset to achieve accurate engagement with the object.
@@
The precision and robustness of this system allow for successful detection and manipulation even with smaller ArUco tags, as demonstrated in the fine alignment tool deployment. To facilitate consistent image acquisition, the workspace is illuminated with diffuse white-light LED strips integrated into the overhead structure. Additional directional LEDs are positioned in the fine-alignment tool to enhance tag visibility. To suppress glare and reduce errors from ambient light sources, light-blocking panels are installed around the frame perimeter.
@@
A potential limitation of the current implementation is the need for uniquely assigned ArUco tags across all components. In practice, this is mitigated by the large dictionary size and spatial separation of simultaneously visible tags. In multi-robot configurations, each robot can utilize a dedicated camera and ArUco ID space, supporting scalability for distributed collaboration in large or reconfigurable optical assemblies.
X.6 Optical component casing and bases
To facilitate reliable robotic handling of diverse optical components, we designed and fabricated universal housings and bases using fused deposition modeling (FDM) 3D printing with polylactic acid (PLA) filament using the Bambu X1E 3D printer. Each housing is tailored to securely enclose optical components, while presenting a standardized external geometry. This uniformity eliminates the need for the robotic arm to adapt dynamically to variations in component shapes and sizes, simplifying grasping and manipulation.
An indentation feature is incorporated into each housing, precisely aligned to the robot’s claw gripper dimensions. This design ensures repeatable, stable engagement of the gripper during pickup and transport, significantly reducing the likelihood of slippage or misalignment during handling.
For optical component positioning and identification, an ArUco fiducial markers are affixed to the top and side surfaces of each housing. Using these fiducial markers, the system achieves robust visual tracking, even under moderate occlusion or variable lighting conditions.
To ensure stable and repeatable placement of optical components, each unit is supported by a custom-designed base optimized for tabletop experimental environments. The bases are fabricated using fused deposition modeling (FDM) 3D printing with polylactic acid (PLA) filament. To prevent sliding and enhance frictional stability, each base is equipped with a rubberized bottom layer. Additionally, multiple neodymium magnets are embedded within the lower section of the base. The combined effect of the rubber and magnets significantly increases the frictional force between the base and the optical table, preventing unwanted sliding, tipping, and rotation under typical laboratory vibrations or minor external disturbances (e.g., those caused by the insertion of the fine-alignment tool).
X.7 Robotic pick-and-place operations
@@
Following coarse localization, the robotic arm approaches the component using a pre-defined trajectory optimized to minimize collision risks. During the final approach phase, a secondary camera mounted on the robot’s end-effector provides high-resolution, close-range visual feedback. This enables fine alignment of the gripper with the standardized indentation designed into each casing, ensuring a secure and precise engagement during pickup.
@@
Before finalizing placement, the end-effector camera performs a local verification and correction routine. Minor translational or rotational deviations are adjusted in situ to ensure alignment within the specified tolerance bounds. This dual-stage visual feedback—combining global position estimation with local fine correction—enables highly accurate, repeatable pick-and-place performance, which is critical for the assembly and realignment of complex optical experimental setups.
X.8 Correction of positional errors
Despite the high accuracy of the robotic placement system, occasional positional errors are observed. These inaccuracies arise from two primary sources: variable magnetic attraction between component bases and the optical table, and mechanical slippage at the gripper interface due to wear of the rubber contact pads.
To mitigate these errors and ensure reliable and robust assembly of setups, we implement an automated correction protocol based on post-placement verification. After each placement, the computer vision system re-evaluates the actual position of the optical element and compares it against the target coordinates. If the distance between the detected and intended positions exceeds a predefined tolerance threshold, the robot initiates a re-placement procedure. For all experiments reported in this work, we set this tolerance threshold to to 0.5 mm.
@@ X.8 Correction of positional errors
X.9 Fine alignment tool
@@
The core actuation mechanism consists of two custom-fabricated Allen keys that serve as precision interface tools for engaging standard optical component knobs. These Allen keys are machined from solid stock with a seamless transition between the shaft and tool head, forming a single integrated axle. The design was iteratively refined through multiple fabrication cycles, achieving sub-100 µm tolerances. To minimize mechanical resistance and reduce power draw, the rotational axes are supported by high-performance bearings and shaft collars, ensuring stable alignment and preventing stalling during actuation.
@@
X.10 Optical measurements
For the fabrication of the multilayer photonic crystals (measured in Figure 4d), we employ plasma-enhanced chemical vapor deposition (PECVD) using the SAMCO PD-220 system to deposit a stack of silicon (Si) and silica (SiO2) layers onto a fused silica substrate (microscope coverslips of thickness 180 microns) with a deposition rate of approximately 55 nm/min for both materials. The deposition time is used to control the thicknesses of the individual layers. The photonic crystal consists of eight total layers, alternating between Si and SiO2, each with a thickness of 40 nm except the fifth (Si) layer of thickness 110 nm which serves as a defect layer, creating a cavity. The simulated angle- and wavelength-resolved transmission spectrum is calculated using the Rigorous Coupled Wave Analysis (RCWA) method, as implemented in Stanford Stratified Structure Solver (S4[19]. In the simulations, the material model (n, k) for Si is taken from Ref.[20] and fit to a fourth order polynomial in the wavelength range of 400-700 nm, while the index of SiO2 is assumed to be 1.45.

### v2

@@ X.3 Validation of optical setups and LLM performance
To ensure that optical setups generated by the LLM were accurate and adhered to physical and geometric constraints, we employed a validation procedure that extended the checks used during optical dataset generation. First, we applied a set of rule-based spatial constraints to the model-generated setup: all components must lie within the robot’s 61 cm radius reach, remain outside the 15 cm radius circular footprint of the robot at the origin, maintain non-overlapping placement based on their 7.5 cm × 6.4 cm physical dimensions, and internal beam paths must not intersect the robotic arm’s footprint.
@@
Passing these spatial checks does not guarantee that the laser beam will correctly propagate through the setup, as the component orientations or positions may still be physically inaccurate. To address this, we performed an additional validation step that verified the structural correctness of the generated setup against a known reference. We did this by attempting to identify a similarity transformation that mapped the LLM-generated setup onto one of the predefined reference setups of the same type. This is framed as a geometric alignment task, where the candidate transformation must preserve component identities and maintain approximate spatial relationships within tolerance bounds. If such a transformation exists, the generated setup is deemed valid due to its equivalence to the corresponding reference configuration. This step ensured that the model had not only obeyed spatial constraints but also produced a layout that structurally corresponds to an accurate optical arrangement of components.
The accuracy of all prompting strategies was measured against a newly generated test dataset with 128 samples for each setup, utilizing a generation temperature of $0.5$ to moderate output randomness. The accuracies and user compliance scores are reported in Figure 2 of the main text.
@@ X.4 Generation of robot motion code
In this part of our workflow, we utilized Claude 3.7, an LLM, to transform high-level interferometer design specifications into executable robotic control code. By leveraging pre-defined wrapper functions that discretize the optical assembly task, we systematically translated the design parameters directly into robotic instructions. The process involved providing Claude 3.7 with a list of tuples with component specifications in the format $\mbox{component ID},\mbox{x position},\mbox{y position},\mbox{orientation angle}]$ from the LLM design along with a mapping between component identifiers and their corresponding ArUco markers of the hardware for computer vision recognition. After describing the system’s functional capabilities through four wrapper functions (system initialization, component pickup via visual recognition, precise placement, and robotic arm homing), Claude generated a Python main function that systematically processed the interferometer design. The resulting code initialized the vision and robotic systems, parsed the design data, appropriately filtered non-robotic components, and implemented an assembly sequence that handled component identification, retrieval, precise placement, and safe arm positioning between operations. This procedure achieved the optical geometry specified by the LLM optical design.
X.5 Robotic arm specifications
Figure S2: Pick-and-place repeatability maps. The robot was programmed to repeatedly transfer an object between two target points (A and B), separated by 5 cm, at a fixed orientation for each point. The scatter of points indicates the final placement positions, while arrows denote the measured orientations at each trial.
The robotic arm is a UFACTORY xArm 7 with 7 degrees of freedom, a payload capacity of 3.5 kg, and a reach radius of 61 cm. The manufacturer reports a nominal repeatability of $\pm$0.1 mm; however, under our experimental conditions and measurement pipeline, we determined practical spatial and angular placement precisions of $\pm$0.5 mm and $\pm$0.2, respectively. These values were obtained by repeatedly commanding the robot to transfer an object between two fixed locations separated by 5 cm, and extracting the resulting distributions of object positions and orientations (Figure S2). The observed scatter plots, which we refer to as pick-and-place repeatability maps, highlight the robot’s capacity for consistent handling and placement of objects.
X.6 Computer vision system
@@
The robotic system’s vision pipeline is built around a stereo vision setup composed of two 4K ELP cameras mounted 23 cm apart. Initial calibration was conducted using a standard checkerboard pattern, with images collected at varying positions on the optical workbench. Intrinsic and extrinsic parameters of the stereo pair were obtained using the OpenCV cv2.calibrateCamera and cv2.stereoCalibrate functions. The calibration outputs were validated by comparing the predicted baseline offset to the physical camera separation, confirming sub-millimeter discrepancies. These calibration parameters were stored in a serialized NumPy archive (stereo_calibration2.npz) for use in subsequent 3D position estimations during coarse alignment operations.
@@
To enable robust and accurate object detection, we employed the ArUco fiducial markers. Each optical component is tagged with a unique identifier drawn from a 6×6 binary dictionary comprising over 800 distinguishable codes (see [28] for code). These tags are detected using the cv2.aruco module, which computes the component’s pose via the cv2.solvePnP algorithm using previously derived camera intrinsics. While the x and y positional estimates from ArUco detection were sufficiently accurate for coarse positioning, the z-coordinate was instead inferred from a depth measurement. For this, we deployed a robot-mounted Intel RealSense depth camera equipped with LiDAR. After the ArUco tag was centered in the camera’s field of view, the LiDAR sensor provided a high-accuracy depth estimate, enabling precise z-coordinate acquisition. The robot’s end-effector was then translated vertically from a known camera-to-gripper offset to achieve accurate engagement with the object.
@@
The precision and robustness of this system allowed for successful detection and manipulation even with smaller ArUco tags, as demonstrated in the fine alignment tool deployment. To facilitate consistent image acquisition, the workspace was illuminated with diffuse white-light LED strips integrated into the overhead structure. Additional directional LEDs were positioned in the fine-alignment tool to enhance tag visibility. To suppress glare and reduce errors from ambient light sources, light-blocking panels were installed around the frame perimeter.
@@
A potential limitation of the current implementation is the need for uniquely assigned ArUco tags across all components. In practice, this is mitigated by the large dictionary size and spatial separation of simultaneously visible tags. In multi-robot configurations, each robot can utilize a dedicated camera and ArUco ID space, supporting scalability for distributed collaboration in large or reconfigurable optical assemblies. The code for computer vision is available in [26, 27]
X.7 Optical component casing and bases
To facilitate reliable robotic handling of diverse optical components, we designed and fabricated universal housings and bases using fused deposition modeling (FDM) 3D printing with polylactic acid (PLA) filament using the Bambu X1E 3D printer. Each housing was tailored to securely enclose optical components, while presenting a standardized external geometry. This uniformity eliminates the need for the robotic arm to adapt dynamically to variations in component shapes and sizes, simplifying grasping and manipulation.
An indentation feature was incorporated into each housing, precisely aligned to the robot’s claw gripper dimensions. This design ensures repeatable, stable engagement of the gripper during pickup and transport, significantly reducing the likelihood of slippage or misalignment during handling.
For optical component positioning and identification, ArUco fiducial markers were affixed to the top and side surfaces of each housing. Using these fiducial markers, the system achieves robust visual tracking, even under moderate occlusion or variable lighting conditions.
To ensure stable and repeatable placement of optical components, each unit is supported by a custom-designed base optimized for tabletop experimental environments. The bases were fabricated using fused deposition modeling (FDM) 3D printing with polylactic acid (PLA) filament. To prevent sliding and enhance frictional stability, each base was equipped with a rubberized bottom layer. Additionally, multiple neodymium magnets were embedded within the lower section of the base. The combined effect of the rubber and magnets significantly increases the frictional force between the base and the optical table, reducing unwanted sliding, tipping, and rotation under typical laboratory vibrations or minor external disturbances (e.g., those caused by the insertion of the fine-alignment tool).
X.8 Robotic pick-and-place operations
@@
Following coarse localization, the robotic arm approaches the component using a pre-defined trajectory that first elevates the end effector above the tallest components in the workspace, thereby minimizing collision risks before descending to engage the target. During the final approach phase, a secondary camera mounted on the robot’s end-effector provides high-resolution, close-range visual feedback. This enables fine alignment of the gripper with the standardized indentation designed into each casing, ensuring a secure and precise engagement during pickup.
@@
Before finalizing placement, the end-effector camera performs a local verification and correction routine. Minor translational or rotational deviations are adjusted in situ to ensure alignment within the specified tolerance bounds. This dual-stage visual feedback—combining global position estimation with local fine correction—enables highly accurate, repeatable pick-and-place performance, which is critical for the assembly and realignment of complex optical experimental setups. The code for robotic arm control is available in [26].
X.9 Correction of positional errors
Despite the high accuracy of the robotic placement system, occasional positional errors were observed. These inaccuracies arise from two primary sources: variable magnetic attraction between component bases and the optical table, and mechanical slippage at the gripper interface due to wear of the rubber contact pads.
To mitigate these errors and ensure reliable and robust assembly of setups, we implemented an automated correction protocol based on post-placement verification. After each placement, the computer vision system re-evaluates the actual position of the optical element and compares it against the target coordinates. If the distance between the detected and intended positions exceeds a predefined tolerance threshold, the robot initiates a re-placement procedure. For all experiments reported in this work, we set this tolerance threshold to 0.5 mm.
@@ X.9 Correction of positional errors
X.10 Fine alignment tool
@@
The core actuation mechanism consists of two custom-fabricated Allen keys that serve as precision interface tools for engaging standard optical component knobs. These Allen keys are machined from solid stock with a seamless transition between the shaft and tool head, forming a single integrated axle. The design was iteratively refined through multiple fabrication cycles, achieving sub-100 µm tolerances. To minimize mechanical resistance and reduce power consumption, the rotational axes are supported by high-performance bearings and shaft collars, ensuring stable alignment and preventing stalling during actuation.
@@
X.11 Optical measurements
For the fabrication of the multilayer photonic crystals (measured in Figure 4d), we employed plasma-enhanced chemical vapor deposition (PECVD) using the SAMCO PD-220 system to deposit a stack of silicon (Si) and silica (SiO2) layers onto a fused silica substrate (microscope coverslips of thickness 180 microns) with a deposition rate of approximately 55 nm/min for both materials. The deposition time was used to control the thicknesses of the individual layers. The photonic crystal consists of eight total layers, alternating between Si and SiO2, each with a thickness of 40 nm except the fifth (Si) layer of thickness 110 nm which serves as a defect layer, creating a cavity. The simulated angle- and wavelength-resolved transmission spectrum was calculated using the Rigorous Coupled Wave Analysis (RCWA) method, as implemented in Stanford Stratified Structure Solver (S4) [29]. In the simulations, the material model (n, k) for Si was taken from Ref.[30] and fit to a fourth order polynomial in the wavelength range of 400-700 nm, while the index of SiO2 was assumed to be 1.45.